# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

### v2

@@ Abstract
Infrared small target detection poses unique challenges due to the scarcity of intrinsic target features and the abundance of similar background distractors. We argue that background semantics play a pivotal role in distinguishing visually similar objects for this task. To address this, we introduce a new task—-clustered infrared small target detection, and present DenseSIRST, a novel benchmark dataset that provides per-pixel semantic annotations for background regions, enabling the transition from sparse to dense target detection. Leveraging this dataset, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Furthermore, we propose the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which selectively pastes small targets into sky regions during training, avoiding the creation of false alarm targets in complex non-sky backgrounds. Extensive experiments validate the effectiveness of BAG-CP and BAFE-Net in improving target detection accuracy while reducing false alarms. The DenseSIRST dataset, code, and trained models are available at https://github.com/GrokCV/BAFE-Net.
Index Terms:
Infrared Small Target Detection; Clustered Target Detection; Dataset; Feature Exchange; Semantic Segmentation
@@ I Introduction
Figure 1: Illustration of the background semantics’ crucial role in infrared small target detection. The small target in the orange box represents a genuine target of interest against the sky background, while the small target in the blue box is a false alarm in the context of the building background. When considering only local regions, it is challenging to differentiate between clustered small targets and false alarms due to their similar appearances. By incorporating global contextual information, the distinction between real targets and false alarms becomes more apparent.
Infrared imaging has emerged as a crucial tool in various domains, particularly in scenarios where long-distance imaging is required in the absence of visible light, such as night-time operations or environments with limited visibility [1]. By capturing the thermal distribution of objects, infrared imaging provides critical information for applications such as unmanned system monitoring, maritime target search [2], and rescue missions [3]. However, in these applications, the targets of interest are often located at great distances, resulting in their appearance as extremely small-sized objects in the captured images. Consequently, infrared small target detection plays a pivotal role in accurately interpreting and extracting valuable insights from these images.
I-A Prior Works on Clustered Infrared Small Target Detection
While the potential of infrared imaging has been recognized across various domains, existing infrared small target detection techniques are predominantly tailored for the detection of sparse targets [4, 5], which proves to be inadequate when faced with complex application scenarios such as drone swarm defense, multi-target tracking systems, and extensive maritime search and rescue operations. Consequently, there is an urgent need for advanced technology capable of effectively identifying and tracking clustered infrared small targets.
As the distribution of small targets transitions from sparse to dense scenarios, existing infrared small target detection techniques face a multitude of challenges that drastically hinder their effectiveness. Firstly, methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies. Moreover, the reliance on local contrast features [6], a staple of traditional approaches, becomes unreliable as the targets are no longer set against a uniform background but are instead interspersed with other proximate objects. Furthermore, even state-of-the-art deep learning methods [7, 8, 9, 10] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, significantly adding to the complexity of detection tasks. For instance, Dai et al. [11] introduced a detection transformer designed for arbitrarily oriented objects, yet its applicability to dense infrared small target settings remains limited due to the lack of intrinsic target features. Therefore, current methods are plagued by inefficiencies when operating in dense settings, frequently resulting in missed detections and high false alarm rates. This calls for new solutions tailored to overcome the intricacies of dense infrared small target detection.
While the rapid advancements in generic object detection have led to state-of-the-art methods like TOOD [12] and CrossKD [13] that address dense object detection, these techniques cannot be directly applied to the unique challenges posed by infrared small target detection. The specialized nature of this field presents several critical factors that hinder the direct transplantation of generic vision approaches:
@@ I-A Prior Works on Clustered Infrared Small Target Detection
    Scarcity of Intrinsic Target Features: The foremost challenge stems from the paucity of intrinsic target features. Unlike generic objects, which boast rich visual cues such as appearance, texture, and structure, infrared small targets are minuscule in size, merely popping out as nondescript blobs in the image. This “visual poverty” renders recognition and localization methods based on target-specific characteristics largely ineffectual.
@@
    Abundance of Similar Background Distractors: Infrared imagery is often plagued by a multitude of background distractors that bear a striking resemblance to the targets of interest. These distractors, such as noise spots and hot spots, also manifest as blobs, making it nearly impossible to discriminate between them and genuine targets based on local visual comparisons. Liao et al. [14] tackled similar challenges in RGB-thermal salient object detection, where distortions like noise and motion blur compounded the complexity of separating salient objects from background clutter.
@@
I-B Motivation
In this paper, we posit that the answer to the aforementioned question lies in the background semantics. As illustrated in Fig. 1, visually similar small objects at a distance can be classified differently based on their surrounding context. A small object in the sky background is likely to be a target of interest, while the same object against a building backdrop might be a false alarm. This stark contrast underscores the critical role that background semantics play in assigning meaningful labels to visually similar objects.
@@ I-B Motivation
However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research. Despite the availability of several public datasets [15, 3, 16, 10], none of these datasets provide detailed semantic annotations for the background regions. To enable the explicit semantic modeling of backgrounds, a prerequisite is the creation of a dataset that includes background class annotations. Such a dataset would provide a solid foundation for researchers to develop algorithms that can leverage the rich contextual information present in the background regions.
To address the aforementioned challenges, this study presents DenseSIRST, a new open-source infrared target detection benchmark dataset. Compared to previous datasets in this field, DenseSIRST has implemented two significant upgrades. Firstly, the dataset transitions from a simple sparse target detection task to the precise detection of dense, clustered targets, substantially increasing the task’s difficulty and practical value. Secondly, DenseSIRST performs detailed per-pixel semantic segmentation of the background, dividing it into sky and non-sky categories, thereby providing preliminary support for explicit semantic modeling of the background.
@@
Building upon this foundation, we propose the Background-Aware Feature Exchange Network (BAFE-Net) framework, which transforms the infrared small target detection paradigm from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. BAFE-Net introduces a dynamic cross-task feature hard-exchange module in the target detection head and background segmentation head, which selectively embeds target semantics and background semantics between the segmentation and detection tasks through a feature adapter. To the best of our knowledge, this paper is the first to propose a cross-task feature hard-exchange mechanism. Benefiting from the explicit modeling of background semantics, BAFE-Net can better understand the contextual information in the scene, thereby significantly improving the accuracy of target detection and greatly reducing the false alarm rate.
Finally, by leveraging the fine-grained semantic annotations of the background, we propose a Background-Aware Gaussian Copy-Paste (BAG-CP) method. During the training phase, this method selectively pastes small targets into the sky region based on different background environments during data augmentation. This approach avoids pasting small targets into complex non-sky background regions, such as urban ground, which can produce highly similar false alarm targets.
@@
    We propose a new task—-clustered infrared small target detection, and have released a new dataset, DenseSIRST, annotated with background semantic segmentation.
@@
    Utilizing DenseSIRST, we propose and demonstrate the effectiveness of the Background-Aware Gaussian Copy-Paste (BAG-CP) method, which outperforms the original Copy-Paste method that ignores background semantics.
@@
II Related Work
II-A Infrared Small Target Detection
Infrared small target detection methods can be broadly classified into two main categories: model-driven and data-driven approaches. The model-driven methods, in particular, encompass three distinct subclasses, each employing unique strategies to capture and represent the relationship between targets and their surrounding backgrounds.
  1. 1.
    Background Estimation Methods: By leveraging the local consistency of backgrounds, these methods construct filters to enhance target saliency. However, they falter when dealing with extremely weak targets, leading to missed detections and a high sensitivity to hyperparameters. The limitation stems from insufficient semantic understanding of the features [17], which results in poor performance in complex backgrounds and increased false alarms.
  2. 2.
    Human Visual System (HVS) Methods: Techniques such as local contrast measure (LCM) [6] and multi-scale patch-based contrast measure (MPCM) [18] attempt to emulate the high spatial local contrast sensitivity of the human eye. Despite their intuitive appeal, these methods struggle with very small or blurred targets, mainly due to their inability to generate sufficient contrast.
  3. 3.
    Low-Rank and Sparse Decomposition (LRSD) Methods: This category includes approaches like the infrared patch-image (IPI) model [4] and reweighted infrared patch-tensor (RIPT) model [5]. These methods have gained popularity due to their superior background suppression capabilities compared to previous techniques. However, they severely suffer from the challenge of distinguishing background distractors from targets due to the similar sparse characteristic.
Despite their contributions, these methods suffer from several limitations, including poor performance in complex backgrounds, high false alarm rates, and sensitivity to hyperparameters, which we attribute to the lack of feature semantics.
Transitioning from model-driven methods, data-driven approaches have revolutionized the field by leveraging deep networks to automatically learn features relevant for detecting small targets. Recent advancements have also focused on implementing these deep learning-based methods efficiently, such as the work on high-accuracy and real-time two-stage small object detection on FPGA [19]. Methods such as asymmetric context modulation (ACM) [15], dense nested attention networks (DNANet) [16], and gated shaped trans U-Net (GSTUnet) [9] have shown significant promise. These methods concentrate on integrating the features of infrared small targets with neural networks, utilizing deeper neural networks or more robust attention modules. Similar advancements in applying deep learning to infrared imaging have been seen in related fields, such as the work on infrared-guided eye localization for autostereoscopic displays [20]. The advantages of data-driven methods lie in their ability to automatically extract features and focus on multi-scale feature interaction, which can be seen as a form of implicit modeling of the relationship between targets and background. Similar approaches have been applied in other domains, such as cloud detection, where multi-level information fusion with edge information injection has been used to improve detection performance [21]. In a related field, transformer-based models have been employed for spatio-temporal unsupervised traffic anomaly detection in aerial videos [22], demonstrating the versatility of these advanced deep learning techniques in handling complex visual data with small objects or anomalies.
While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances. Moreover, despite not fully depending on sparse priors or local contrast priors like traditional methods, some approaches still leverage this prior knowledge to enhance detection performance [23]. Their performance is greatly affected in dense scenes with multiple adjacent targets, where the assumption of local contrast between targets and their neighboring background no longer holds. This fundamental shift in target distribution necessitates the development of approaches that can effectively handle the complex interactions and dependencies among multiple targets in close proximity.
In contrast to previous studies, our work explores a new detection paradigm for infrared small targets:
  1. 1.
    Unexplored Challenge for Clustered Targets: Unlike traditional approaches that focus on sparse targets, we take on the more complex task of detecting clustered targets, introducing a new challenge in this field. We further facilitate the research by providing DenseSIRST, a dedicated dataset created specifically for this novel task.
  2. 2.
    Explicit Semantic Modeling: In contrast to previous studies that treat backgrounds as a homogeneous class, we leverage the DenseSIRST dataset’s semantic annotations for background modeling. Our proposed BAFE-Net thereby gains an enhanced understanding of the target’s background, leading to improved detection accuracy.
  3. 3.
    Introduction of Cross-Task Feature Exchange: Our dynamic cross-task feature exchange module brings together background semantics and target features through a feature adapter, a strategy not found in existing work. This approach enables a more precise differentiation between targets and their backgrounds.
II-B Data Augmentation for Object Detection
Data augmentation is pivotal in enhancing the performance and generalization of object detection models. By diversifying training data through various transformations such as flipping, rotation, scaling, and cropping, models learn to recognize and generalize across a broader spectrum of object variations. However, while these basic techniques are widely used, they often fall short in complex detection scenarios, necessitating more sophisticated methods.
To address the limitations of traditional methods, advanced augmentation techniques have been developed, which introduce more nuanced transformations that better preserve and highlight essential features of the objects while still enhancing the dataset’s diversity. For instance, CutMix [24] improves upon the CutOut [25] method by overlaying a segment of one image onto another, blending labels proportionally to the area of the patch, thereby maintaining more contextual integrity. Attentive-CutMix [26] refines this approach by selecting regions based on feature importance, thus preserving significant target details. SnapMix [27] and SaliencyMix [28] adjust the blending process based on saliency or importance maps, aiming to preserve crucial features while enhancing background complexity. TransMix [29] integrates attention mechanisms to better manage the blending of features, focusing on maintaining important target characteristics during augmentation. StyleMix [30] and PuzzleMix [31] introduce style and structural considerations into the mix, allowing for more complex transformations that consider both content and style elements of the images, potentially offering more robust augmentation strategies.
Despite these advancements, these methods primarily assume the presence of larger, easily distinguishable targets and do not specifically cater to the unique challenges posed by clustered infrared small target detection. None of these methods are explicitly designed to handle the high-density and occlusion-prone nature of infrared small targets. This gap highlights the need for augmentation strategies that directly address the intricacies of detecting small, closely-spaced targets against complex backgrounds.
Although our BAG-CP aligns with the broader goal of enhancing infrared small target detection, it stands out in its unique focus and strategies:
  1. 1.
    Pioneering Data Augmentation for Clustered Infrared Small Target Detection: Uniquely, our method pioneers a data augmentation technique specifically tailored for clustered infrared small target detection. It takes into account the local contrast prior and enables dynamic random variations in local contrast, a perspective not considered in existing techniques.
  2. 2.
    Background-Aware Target Placement: Our BAG-CP approach diverges from the commonly used random or fixed-position copy-paste tactics. Instead, it integrates background semantics to guide target placement, thereby skirting potential category conflicts induced by inconsistent background contextual semantics.
II-C Multi-Task Learning and Cross-Task Mechanism
Multi-task learning (MTL) is a pivotal strategy in computer vision that enhances model performance and generalization by jointly learning shared representations from related tasks, such as detection and segmentation. Takikawa et al. [32] introduced a two-stream CNN architecture for semantic segmentation that explicitly wires shape information as a separate processing branch, using gates to connect the intermediate layers of the two streams. Misra et al. [33] proposed a new sharing unit called “cross-stitch” unit, which combines activations from multiple networks and can be trained end-to-end to learn an optimal combination of shared and task-specific representations. Zhang et al. [10] developed ISNet, which employs a Taylor finite difference inspired edge block and two-orientation attention aggregation block to enhance edge information and capture shape characteristics of infrared small targets.
However, these methods may not be well-suited for clustered infrared small target detection, as they often rely on edge detection or surface normal estimation, which assume larger targets. In real-world scenarios, infrared small targets typically consist of only 1 to 9 pixels and are subject to the diffusion effect, lacking clear edges and falling far below the conventional definitions of small targets, i.e., $32\times 32$ pixels in general object detection.
In this context, cross-task mechanisms that enable semantic exchange between different tasks are more relevant to our work. These methods enhance MTL by facilitating feature exchange or distillation across different tasks, allowing each task to benefit from complementary signals of others. Methods like prediction-and-distillation network utilize attention-based mechanisms to distill multi-modal features [34], while MTI-Net extends this concept with a multi-scale approach better suited for handling diverse task demands [35].
While our research shares the underpinnings of multi-task learning and cross-task mechanisms, we pioneer a new approach to feature fusion. Instead of following the well-trodden path of soft fusion found in attention-based methods, our work introduces an innovative protocol–-a dynamic hard-exchange mechanism. To the best of our knowledge, this is the first work to explore the potential of dynamic hard-exchange of features between tasks.
III The DenseSIRST Dataset
Figure 2: Illustration of selected images from our proposed DenseSIRST dataset. Each image pair consists of a simulated dense small target image (top) and its corresponding sky-segmented version (bottom). These images highlight the challenges posed by densely clustered targets within diverse backgrounds.
III-A Motivation and Rationale
The detection of small, low-contrast targets against complex backgrounds in infrared imagery poses a significant challenge. Existing datasets, though instrumental in progressing the field, exhibit critical deficiencies. As shown in Table I, well-regarded datasets such as SIRST V1 [15], SIRST V2 [3], IRSTD1K [10], and SIRSTAUG [36] predominantly feature sparsely distributed targets and lack per-pixel background annotations. This stands in stark contrast to the dense, diminutive targets frequently encountered in real-world scenarios.
To bridge this gap, we present DenseSIRST: a meticulously annotated dataset of infrared images teeming with densely clustered, small targets, which is a more realistic and challenging benchmark for algorithm training and evaluation. Fig. 2 showcases a selection of images from the DenseSIRST dataset, illustrating the dense distribution of small targets and the corresponding sky segmentation annotations. These annotations provide valuable contextual information for developing and evaluating detection algorithms.
@@ III-A Motivation and Rationale
TABLE I: Comparison of the DenseSIRST dataset with other popular datasets used in infrared small target detection.

### v3

@@ Abstract
Infrared small target detection presents significant challenges due to the limited intrinsic features of the target and the overwhelming presence of visually similar background distractors. We contend that background semantics are critical for distinguishing between objects that appear visually similar in this context. To address this challenge, we propose a task—clustered infrared small target detection—and introduce DenseSIRST, a benchmark dataset that provides per-pixel semantic annotations for background regions. This dataset facilitates the shift from sparse to dense target detection. Building on this resource, we propose the Background-Aware Feature Exchange Network (BAFE-Net), a multi-task architecture that jointly tackles target detection and background semantic segmentation. BAFE-Net incorporates a dynamic cross-task feature hard-exchange mechanism, enabling the effective exchange of target and background semantics between the two tasks. Comprehensive experiments demonstrate that BAFE-Net significantly enhances target detection accuracy while mitigating false alarms. The DenseSIRST dataset, along with the code and trained models, is publicly available at https://github.com/GrokCV/BAFE-Net.
@@ I Introduction
Figure 1: Illustration of the background semantics’ crucial role in infrared small target detection. When considering only local regions (as in the top-left image), it is challenging to differentiate between clustered small targets and false alarms due to their similar appearances. By incorporating global contextual information (as in the surrounding images), the distinction between real targets and false alarms becomes more apparent. From the global images, it can be seen that the sky background is highlighted in red. The small target in the blue box represents a genuine target of interest against the sky background, while the small target in the red box is a false alarm in the context of the building background.
Infrared small target detection has emerged as a fundamental challenge in computer vision, with critical applications spanning military surveillance, drone detection, search and rescue missions, and night operations. The task involves identifying and localizing diminutive objects that appear as few pixels in infrared imagery, often captured under adverse weather conditions or limited visibility scenarios. As defense technologies evolve and multi-target scenarios become increasingly prevalent, the demand for robust and accurate detection systems has intensified dramatically.
I-A Challenges on Clustered Infrared Small Target Detection
Traditional infrared small target detection research has predominantly focused on sparse target scenarios [1, 2], where isolated objects are distributed across relatively uniform backgrounds. However, contemporary real-world applications present fundamentally different challenges. In defense against drone swarms, multi-target tracking systems, and large-scale maritime rescue operations, targets frequently appear in clustered configurations, creating dense distributions that fundamentally alter the detection paradigm. This gap underscores an urgent need for advanced technology that can effectively identify and track clustered infrared small targets.
This transition from sparse to dense target detection introduces a cascade of technical challenges that render existing methodologies inadequate:
@@ I-A Challenges on Clustered Infrared Small Target Detection
    Methods that rely on the assumption of sparsely distributed targets become inapplicable in dense scenes, where targets are closely situated and may occlude one another, leading to the breakdown of single-target analysis strategies.
@@
    Local contrast based methods [3], which depend on detecting objects against uniform backgrounds, become unreliable as dense clusters of targets blur the line between target and background.
  3. 3.
    Even state-of-the-art deep learning methods [4, 5, 6] struggle when confronted with dense target scenarios where the delineation between targets and background blurs, adding to the complexity of detection tasks.
These challenges are compounded by the limitations in current datasets, which primarily focus on sparse target detection and fail to address the complexities of clustered targets.
I-B Limitations of Current Approaches
While the rapid advancements in generic object detection have led to state-of-the-art methods like TOOD [7] and CrossKD [8] that address dense object detection, these techniques cannot be directly applied to the unique challenges posed by infrared small target detection. The specialized nature of this field presents two critical factors that hinder the direct transp lantation of generic vision approaches:
  1. 1.
    Scarcity of Target-Specific Features: Infrared small targets typically occupy only a few pixels and lack distinctive visual characteristics such as texture, shape, or structural details. This “visual poverty” renders traditional feature based recognition and localization methods largely ineffective, as there are insufficient discriminative cues to differentiate targets from background elements.
  2. 2.
    Abundance of Similar Background Distractors: Infrared imagery frequently contains background elements [9], such as thermal noise, hot spots, and structural artifacts, that exhibit visual similarity to genuine targets. This similarity creates a high false alarm rate when relying solely on local appearance-based discrimination.
@@ I-B Limitations of Current Approaches
I-C Motivation
The answer to the aforementioned challenge lies in a paradigm shift: rather than relying solely on target appearance, we propose to exploit background semantics.
Consider the scenario illustrated in Fig. 1: when examining only local regions, distinguishing between genuine targets and false alarms becomes nearly impossible due to their visual similarity. However, by incorporating global contextual information, the classification task becomes significantly more tractable. A small bright object appearing against a sky background typically represents a target of interest, while an identical object against building infrastructure is more likely a false alarm. This contextual disambiguation demonstrates that background semantics[10] serve as crucial semantic anchors that enable accurate target identification where local appearance fails.
@@ I-C Motivation
However, it is crucial to highlight a significant gap in the current landscape of infrared small target detection research. Despite the availability of several public datasets [11, 12, 13, 5], none of these datasets provide detailed semantic annotations for the background regions.
Without these annotations, it is challenging to explicitly model the background semantics necessary for improving detection performance. To bridge this critical gap and validate our hypothesis, we introduce DenseSIRST. This is a comprehensive infrared target detection benchmark that addresses both the clustered target detection challenge and the background semantic modeling requirement. Our dataset implements two revolutionary improvements over existing resources:
  • Dense Target Scenarios: Transitioning from sparse to clustered target configurations that reflect real-world deployment conditions such as drone swarm detection and multi-target tracking applications.
  • Background Semantic Annotations: Providing detailed per-pixel semantic segmentation of background regions, categorized into sky and non-sky classes, thereby enabling explicit background semantic modeling for the first time in this domain.
@@
Building upon the semantic-rich DenseSIRST dataset, we propose the Background-Aware Feature Exchange Network(BAFE-Net) framework, which fundamentally transforms infrared small target detection from a single target detection task that solely focuses on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. The core innovation of BAFE-Net lies in its dynamic cross-task feature hard-exchange module, which facilitates selective semantic information transfer between detection and segmentation tasks through an adaptive feature adapter. This represents the first implementation of cross-task feature hard-exchange mechanisms in infrared target detection. By explicitly modeling background semantics, BAFE-Net achieves superior contextual understanding, resulting in significant improvements in detection accuracy while substantially reducing false alarm rates.
@@
    We propose a new task—-clustered infrared small target detection, which poses unique challenges compared to traditional single-target detection.
@@
    We have released a new dataset, DenseSIRST, annotated with background semantic segmentation. This dataset offers rich contextual information to enhance target detection.
@@
TABLE I: A Comparative Analysis of the DenseSIRST Dataset and Other Popular Datasets for Infrared Small Target Detection.
@@
II Related Work
The core idea of traditional infrared small target detection algorithms is to suppress background interference and enhance target saliency. These methods mainly fall into three categories:
  1. 1.
    Background Estimation Methods: By exploiting the local smoothness or non-local self-similarity of background regions, various filtering strategies are designed to highlight targets. However, these methods often fail when facing extremely weak or low-contrast targets, primarily due to their limited semantic perception capabilities and high sensitivity to handcrafted hyperparameters. Consequently, they are prone to both missed detections and false alarms in complex backgrounds.
  2. 2.
    Human Visual System (HVS) Methods: Techniques such as local contrast measure (LCM) [3] and multi-scale patch-based contrast measure (MPCM) [15] attempt to emulate the high spatial local contrast sensitivity of the human eye. While intuitive and computationally efficient, their reliance on pixel-level or patch-level contrast leads to diminished performance when the target is extremely small, blurred, or embedded in cluttered textures.
  3. 3.
    Low-Rank and Sparse Decomposition Methods: Inspired by matrix and tensor decomposition theories, methods such as the Infrared Patch-Image (IPI) model [1], the Reweighted Infrared Patch-Tensor (RIPT) model [2] and the low-rank tensor approximation with saliency prior (LRTA-SP) model [16] decompose the image into low-rank background and sparse target components. Despite demonstrating strong background suppression capabilities, they struggle to differentiate true targets from structured background noise or distractors due to the shared sparsity characteristics, limiting their robustness in real-world scenarios.
With the advent of deep learning, data-driven methods have significantly advanced infrared small target detection by learning robust representations directly from data, rather than relying on handcrafted priors. These approaches leverage deep neural networks to automatically extract semantic features, enabling better adaptation to diverse and complex backgrounds.
Several representative methods have demonstrated the advantages of deep architectures in modeling contextual information and capturing multi-scale cues. For example, the asymmetric context modulation network (ACM) [11] introduces a context-sensitive structure to enhance the discrimination between targets and background. DNANet [13] further incorporates dense attention modules to refine features hierarchically, boosting detection performance across varied scenes.
Recently, FATCNet [17] employs a dual-branch CNN-Transformer with adaptive fusion for effective local-global feature integration. Similarly, ILNet [18] enhances detection by emphasizing low-level salient features via lightweight fusion and dynamic aggregation. Moreover, MNHU [19] introduces a high-order UNet with nested multi-scale fusion to capture long-range dependencies and hierarchical representations. In addition, SFCANet [20] combines spatial-frequency attention with deep supervised multi-task learning to improve background suppression and target discrimination.
While effective to a degree, these methods treat all backgrounds as a single class, ignoring the significant impact of different background semantics on distinguishing between true targets and false alarms with similar appearances. This simplification limits the ability of the model to exploit contextual cues embedded in semantically distinct regions such as sky, sea, or land, which often exhibit different noise characteristics and target distributions.
Moreover, although deep learning-based methods no longer explicitly rely on sparse priors or local contrast assumptions as traditional approaches do, many still implicitly incorporate such prior knowledge to improve detection performance. However, in dense scenes characterized by multiple closely spaced targets, the assumption of local contrast between a target and its surrounding background becomes invalid. This shift in spatial distribution not only weakens the effectiveness of contrast-based cues but also introduces significant ambiguity in distinguishing adjacent targets. Consequently, there is a pressing need for methods capable of modeling the complex interactions and contextual dependencies among targets in densely populated scenarios.
To address these limitations, our work explores a novel detection paradigm for infrared small targets, focusing on two main aspects:
  1. 1.
    Unexplored Challenge for Clustered Targets: Departing from conventional assumptions of sparsity, we target the more realistic and challenging scenario of detecting densely clustered targets. To support this direction, we construct DenseSIRST, a specialized dataset designed to benchmark performance under such conditions.
  2. 2.
    Explicit Semantic Modeling: Rather than treating all backgrounds uniformly, we utilize the semantic annotations provided by DenseSIRST to enable fine-grained background-aware detection. Our proposed BAFE-Net leverages this information to enhance contextual understanding, thereby improving the discrimination between targets and background distractors.
III The DenseSIRST Dataset
Figure 2: Synthetic generation pipeline and representative samples of the DenseSIRST dataset. (a) The BAG-CP pipeline for synthesizing the DenseSIRST dataset incorporates three key components: semantic-aware background selection, Gaussian-based target fusion, and realistic sample generation. (b) Illustration of selected images from our proposed DenseSIRST dataset. Each image pair consists of a simulated dense small target image (left) and its corresponding sky-segmented version (right).
Infrared small target detection is crucial in surveillance, early warning, and remote sensing. However, detecting small, low-contrast targets against complex and cluttered backgrounds remains a significant challenge. Existing benchmark datasets, such as SIRST V1 [11], SIRST V2 [12], IRSTD1K [5], and SIRSTAUG [14], have contributed substantially to algorithmic advancement, yet they predominantly feature sparsely distributed targets and lack comprehensive pixel-level background annotations, as demonstrated in Table I.
The mismatch between current benchmarks and real-world scenarios hinders the development of methods for detecting dense, low-contrast targets within complex and semantically rich backgrounds. To address this gap, we present DenseSIRST, a novel infrared dataset tailored for dense infrared small target detection under complex backgrounds. DenseSIRST offers high-density target distributions and fine-grained pixel-level annotations, as illustrated in Fig. 2 (b). By providing richer spatial and semantic context, it aims to foster the development of more robust and practically applicable detection algorithms.
III-A Dataset Construction
The DenseSIRST dataset is constructed by integrating synthetically generated small targets into infrared background images from the publicly available SIRST V2 [12] dataset. To achieve realistic and contextually consistent fusion, we propose a novel strategy called Background-Aware Gaussian Copy-Paste (BAG-CP), which builds upon the principles of mixup [21]. Unlike naive copy paste, BAG-CP selectively pastes targets into semantically regions informed by background context, thereby enhancing the naturalness and variability of synthesized samples. The overall framework is depicted in Fig. 2 (a).
III-A1 Infrared Small Target Library Construction
To simulate diverse and realistic small targets, we first build a target library by extracting individual small target patches from infrared datasets. Each extracted target is carefully cropped to include the target region with minimal background. This target library provides a rich source of varied target appearances, shapes, and sizes for subsequent synthesis.
III-A2 Semantic-Aware Target Paste Region Selection
To ensure plausible and contextually consistent target placement, the BAG-CP strategy incorporates semantic-aware selection of paste regions. Specifically, from each background image from SIRST V2, one to three candidate regions are randomly defined within relatively smooth and homogeneous sky areas. Each selected region is fixed to a size of $20\times 20$ pixels and serves as a dense paste area for simulating clustered target distributions.
III-A3 Target Pasting and Fusion Process
Within each selected region, 6 to 12 targets are randomly sampled from the target library and pasted with controlled scale and spacing. Each target is resized to a dimension no larger than $5\times 5$ pixels to mimic realistic scale variation. Targets are spaced with 1 to 2 pixels apart to emulate natural cluster formations typical in dense small target scenarios. To seamlessly blend the pasted targets with the background, an adaptive Gaussian fusion matrix $G$ is applied, producing smooth transitions at target edges and avoiding harsh visual boundaries. The synthesized target patch $T_{new}$ is computed as:
$T_{new}=B_{raw}+T_{raw}\times\lambda\times G\ ,\lambda\in[0.5,1],$ (1)
where $B_{raw}$ is the original background patch, $T_{raw}$ is the raw target instance, and $\lambda$ is a random brightness scaling factor introducing intensity diversity. The Gaussian matrix $G$ is defined by
where $w$ and $h$ represent the width and height of the target, respectively. The parameters $\rho_{x}$ and $\rho_{y}$ specify the Gaussian center offset relative to the target dimensions, randomly sampled from $[0,0.2]$. The standard deviations $\sigma_{x}$ and $\sigma_{y}$, controlling the spread of the Gaussian kernel, are independently sampled from $[0.3,0.6]$, which enables varied edge softness and thus improves the naturalness of target boundaries.
III-A4 Geometric and Appearance Augmentation
To further increase target diversity, each pasted target undergoes a random rotation transformation with angle $\theta$ uniformly sampled from $90^{\circ},90^{\circ}]$, implemented by
where $(\mu_{x},\mu_{y})$ is the center of the target patch.